# Skill Discovery for Software Scripting Automation via Offline Simulations with LLMs

### v1

@@ Abstract
Skill Discovery for Software Scripting Automation via Offline Simulations with LLMs
@@ 2.2 Program Synthesis with LLMs
Program synthesis aims to generate code given a natural language description Zhang et al. (2023), where LLMs Austin et al. (2021); Chen et al. (2021); Nijkamp et al. (2023) have demonstrated impressive performance. Modern code generation models are typically evaluated on functional correctness Liu et al. (2024b), often requiring predefined unit tests Chen et al. (2021); Li et al. (2022). In the context of software scripting automation, execution results from the software environment provide immediate feedback on generated code. Unlike the setting of code generation with given task descriptions, this study emphasizes deciding what tasks to generate in specific software, leveraging LLMs’ strong code generation capacity that can be further improved with execution feedback from the software environment during offline simulation Kim et al. (2024); Pan et al. (2024).
@@ 2.3 Automation with LLM Agents
With the rapid advancement of Large Language Models (LLMs), researchers have developed systems and benchmarks to automate complex tasks requiring multiple applications in computer environments Xie et al. (2024); Cao et al. (2024). However, existing LLM-based agents face significant challenges in reliably automating these tasks, achieving only about a $15\%$ success rate across several hundred tasks. Efforts to improve LLM capabilities have also focused on domain-specific tasks, such as spreadsheet manipulation Li et al. (2024); Ma et al. (2024b), Graphical User Interface (GUI) automation Gao et al. (2024), and web browsing Yao et al. (2022); Ma et al. (2024a); Deng et al. (2024). In these specialized domains, LLM-based agents demonstrate higher capacity. Our paper aligns with this body of work by leveraging LLMs to automate domain-specific computer tasks. However, we present the first attempt to (1) automate tasks through the software’s internal scripting environment, and (2) use LLM-based agents to systematically explore and identify which tasks can be automated in the software.
@@ 3.1 Task Creation
We rely on an LLM to generate the tasks. To fully exploit the functionalities supported by the software, we adopted two strategies to guide LLMs to “search” for possible tasks, i.e., a top-down approach that starts from high-level functionalities of the software, followed by a bottom-up approach that considers low-level APIs supported by the software.
@@
A unique opportunity and challenge in the Software scripting interface setting is that we have a list of all APIs supported by the software, denoted as $\mathcal{A}=\{a_{1},a_{2},...,a_{n}\}$, and the scripts largely rely on the “collaboration” of multiple APIs. Therefore, we prompt LLMs with appropriate API combinations to spark LLMs’ knowledge about the software. Given the large number of APIs and the potential for multiple tasks to be accomplished using the same set of APIs, we introduce an API synergy graph, $\mathcal{G}$, to model the likelihood of APIs working together. Graph-based representations are widely employed to model relationships in related applications such as API recommendations Qi et al. (2022); Huang et al. (2022) and LLM-based reasoning Anokhin et al. (2024); Liu et al. (2024a).
@@ 3.2 Skill Generation with Trials
One advantage of generating skills offline is that we can generate the code with multiple trials offline without adding any burden to the users during runtime, providing a better user experience. For each generated $t_{i}$, we use a strategy similar to the ideas of , where the LLM learns from the execution feedback in the software to refine its own outputs. Additionally, we use another LLM or a Large Vision Language Model (LVLM) as a validator to judge the script by looking at the generated code, execution output, and (visual) outcome in the software. The validator comments on whether the skill accomplishes the task and provides feedback for improvement for the next trial. A skill is added to the skillset if it passes the validator.
@@ 4 Experiments
Adobe Illustrator is a leading vector graphics software used by professionals worldwide for tasks such as creating logos, illustrations, and complex design elements. It supports multiple categories of high-level functionalities, including drawing, arranging objects, and applying effects. A full list of the high-level functionalities used in the top-down search in Section 3.1 is provided in Appendix A, which can be obtained through Illustrator’s official scripting guide.  [^3] https://helpx.adobe.com/pdf/illustrator_reference.pdf Modern software platforms that support scripting interfaces offer tasks with varying levels of complexity. This study focuses on generating “atomic” skills: fundamental, modular tasks that involve minimal design choices and serve as building blocks for more complex operations. For instance, in the context of Illustrator, instead of creating an entire flower design, an atomic skill would arrange pre-designed petals into a circular pattern, emphasizing precision and modularity. Illustrator’s scripting interface supports $1818$ API endpoints, where $378$ are the methods and the remaining are attributes of the object, enabling extensive programmatic control over the software.
@@
Adobe Illustrator, as a vector design tool, focuses on creating and manipulating graphical objects. In practical applications, a typical Illustrator project often contains numerous objects, and scripts are usually expected to operate on specific subsets of these objects. The common approach is to manually select the desired objects before running the script, leveraging the selection attribute available for each object. Therefore, for each skill, we prompt the LLM to generate initialization scripts that set up the document with necessary elements and adjust their selection attributes according to the task description. Following previous work with other virtual environments Wang et al. (2024b), we prompt the LLM to make the task code generic and reusable, as well as other Illustrator-specific instructions. The full prompt is shown in Appendix B. Table 1 presents an example skill, including the corresponding task description, code implementation, the layout after running the initialization script, and the layout after running the skill.
We now detail how we operationalize our framework in Adobe Illustrator. We first execute the top-down simulation using the functionality categories listed in Appendix A, where the LLM explores tasks in each subcategory over three rounds. The validated skills, combined with sample scripts provided by the software, are then used to construct the API synergy graph. To train the GCN model for link prediction, we randomly split the sampled edges (both positive and negative) into training, development, and test sets with a ratio of 0.85/0.05/0.1. Next, for each of the approximately 500 method APIs, we retrieve its top-$k$ ($k=5$) synergistic APIs and prompt the LLM to generate related tasks for one round. In each round of the task creation stage (both top-down and bottom-up approaches), the LLM generates ten tasks. For each task, the LLM is allowed up to three trials with previous feedback to generate the code of skills. We note that these hyperparameters may be subject to change based on cost considerations. We chose these values to ensure that the cost and size of the generated skillset remain manageable.
@@
To curate a test set that represents the needs of real users, we iteratively prompt the LLM to generate useful tasks in Illustrator. We manually verify the tasks and correct any issues in the corresponding initialization scripts, if any, to enable automatic evaluation, resulting in $94$ test tasks. We use the all-mpnet-base-v2 model from sentence BERT Reimers and Gurevych (2019) to encode the task descriptions for both test set and the skillset built during offline simulation. We use the test task description as the query and retrieve the most relevant skill from the skillset with semantic matching through cosine similarity.
@@
The baseline method for this task is runtime generation where an LLM generates code to solve the query during runtime. We use gpt-4o 2024-08-06 as the LLM for baseline, as well as the ones in the proposed framework for fair comparisons.
@@ 5.1 Effectiveness in Real-World Scenarios
Our proposed method demonstrates a substantial improvement over the baseline in automating tasks within Adobe Illustrator’s scripting environment. As shown in Table 2, our approach achieves a success rate of $44.7\%$, outperforming the baseline’s success rate of $28.7\%$, on the held-out test set. This evaluation setup effectively measures the end-to-end performance of automation systems when deployed in real-world software contexts, with the natural language task descriptions in the test set simulating real user queries.
A key advantage of our approach lies in its markedly lower runtime costs, which is a critical consideration given the potential volume and repetitive nature of user queries. Our approach demonstrates significantly lower response time by avoiding runtime code generation; instead, it retrieves solutions from a pre-built skillset. Additionally, the proposed method requires zero runtime token cost, as the cost during offline simulation is one-off.
These results highlight the benefits of our method in terms of efficiency and resource utilization. By pre-generating a repository of verified skills, we enable rapid retrieval and execution of automation tasks without the latency and cost associated with real-time LLM interactions.
@@
Ours 44.7% 0.1 s 0*
Table 2: Evaluation of the baseline and the proposed approach. * The token cost for offline simulation in our approach is one-off and not included in the runtime cost.
@@ 5.2 Effectiveness of Two Simulation Strategies
This outcome is expected, given the complexity and long-tail nature of software scripting API design. The long tail represents numerous specialized APIs that are infrequently used and therefore less likely to be well-represented in the LLMs’ training data. This highlights room for future improvement, which we discuss in Section 6.
@@
To better understand the differences in skills between the two simulation strategies, we can use skills related to the functionality of “arrange objects” as an example. The top-down simulation produces skills such as arranging selected objects in circle (Table 1), as well as other common patterns like zigzag or starburst formations. hese skills align with typical design tasks and provide straightforward implementations of the "arrange objects" functionality. Meanwhile, the bottom-up simulation produces skills that utilize more specialized APIs and offer advanced functionalities. For example, it might generate a skill that arranges selected objects based on custom tags assigned to them. This allows for dynamic arrangements driven by metadata, which is expected to be rarer in practical usage but provides greater flexibility for advanced users seeking more control over their design elements.
@@
During bottom-up simulation, we use a GNN-based link prediction model to model the synergy of APIs. We show that this model can find synergistic APIs more effectively, compared to semantically matching similar APIs. We evaluate the performance of the GCN model in retrieving synergistic APIs for task creation using the Hit@$k$ metric. In this context, Hit@$k$ evaluates, for each API, how often the correct items are within the top-$k$ predicted items where correct items refer to APIs in existing skills (not in the training set). We compare the link prediction model with a baseline method-semantic matching. For each API, we retrieve top-$k$ similar APIs based on the semantic representation of their descriptions. Similarly, we use all-mpnet-base-v2 for semantic representation and cosine similarity as the similarity measure. Table 5 shows our link prediction model significantly outperforms the semantic matching method in identifying synergistic APIs, demonstrating its effectiveness in enhancing task creation during the bottom-up simulation.
@@
One key advantage of our offline simulation is the ability to perform multiple trials during code generation without incurring runtime penalties. We analyzed the impact of allowing up to three trials for code generation on the overall success rate. Table 6 demonstrates that the success rate nearly doubles after three trials for both top-down and bottom-up strategies. Specifically, the success rate for top-down skills increases from $16.7\%$ after the first trial to $34.9\%$ after three trials. For bottom-up skills, the success rate rises from $23.1\%$ to $46.6\%$. These results underscore the importance of iterative code generation in enhancing the quality and reliability of the generated scripts.
@@ 5.3 Reliability of LVLM’s judgment
Assessing the success of generated scripts requires reliable verification. We employed an LVLM to evaluate whether the execution outcomes satisfy the task descriptions. To estimate the reliability of the LVLM’s judgments, we compared its assessments with human evaluations on a sample of $122$ tasks.
The confusion matrix in Table 7 shows that the LVLM has a precision of $90.9\%$ and a recall of $80.0\%$. Specifically, out of $50$ tasks deemed successful by humans, the LVLM correctly identified $40$ as successful (true positives) and $10$ as failures (false negatives). Meanwhile, the LVLM only incorrectly labeled $4$ tasks as successful when they were not (false positives). These results indicate that the LVLM is generally reliable, albeit slightly conservative in its judgments.
@@
6 Future Work
This study is one of the first to utilize software scripting interfaces as a testbed for LLM-based systems. This setup offers the advantage of reliable and immediate execution feedback in a controlled environment, which is often challenging to obtain in contexts where API interactions involve multiple external services.
Despite the promising results, there is considerable scope for enhancing our method’s effectiveness. One direction for future work is to enrich LLMs’ knowledge of specific software. Our current skillset does not fully exploit the entire range of available APIs, especially those that are less frequently used or more complex. Training a domain-specific LLM tailored to the software through iterative trial and error could generate more accurate and comprehensive skills, thereby enhancing the system’s overall performance.
Another avenue for future research is leveraging real-world usage data. Our skillset is currently derived from static documentation rather than live usage patterns. Mining API interactions from actual scripting projects could provide valuable insights into common user behaviors and task patterns, making the generated skills more applicable in practice. However, obtaining such data is challenging and raises interesting questions about what kinds of behavioral data are useful for mining such patterns and how to collect them responsibly. Additionally, integrating user feedback loops would allow for continuous refinement of skill discovery and adaptation, ensuring that the system evolves in alignment with user needs.
While our method is tailored for software scripting automation, it has the potential to be applied to other API orchestration scenarios. Extending this approach to different software environments or domains could uncover new challenges and opportunities. Applying our offline simulation framework to other applications could help validate its effectiveness and adaptability, potentially bridging the gap between AI capabilities and user needs in specialized domains.

### v2

@@ Abstract
Skill Discovery for Software Scripting Automation
via Offline Simulations with LLMs
@@ 2.2 Program Synthesis with LLMs
Program synthesis aims to generate code given a natural language description Zhang et al. (2023), where LLMs Austin et al. (2021); Chen et al. (2021); Nijkamp et al. (2023) have demonstrated impressive performance. Modern code generation models are typically evaluated on functional correctness Liu et al. (2024b), often requiring predefined unit tests Chen et al. (2021); Li et al. (2022). For example, pass@k@k Chen et al. (2021); Li et al. (2022) evaluates the model’s chance of passing all unit tests with any of $k$ generated samples. In the context of software scripting automation, execution results from the software environment provide immediate feedback on generated code. Unlike the setting of code generation with given task descriptions, this study emphasizes deciding what tasks to generate in specific software, leveraging LLMs’ strong code generation capacity that can be further improved with execution feedback from the software environment during offline simulation Kim et al. (2024); Pan et al. (2024).
@@ 2.3 Automation with LLM Agents
With the rapid advancement of Large Language Models (LLMs), researchers have developed systems and benchmarks to automate complex tasks requiring multiple applications in computer environments Xie et al. (2024); Cao et al. (2024). However, existing LLM-based agents face significant challenges in reliably automating these tasks, achieving only about a $15\%$ success rate across several hundred tasks. Efforts to improve LLM capabilities have also focused on domain-specific tasks, such as spreadsheet manipulation Li et al. (2024); Ma et al. (2024b), Graphical User Interface (GUI) automation Gao et al. (2024); Nguyen et al. (2025), and web browsing Yao et al. (2022); Ma et al. (2024a); Deng et al. (2024). In these specialized domains, LLM-based agents demonstrate higher capacity. Our paper aligns with this body of work by leveraging LLMs to automate domain-specific computer tasks. However, we present the first attempt to (1) automate tasks through the software’s internal scripting environment, and (2) use LLM-based agents to systematically explore and identify which tasks can be automated in the software. Our primary objective is to generate a verified skillset that represents a software’s supported functionalities and aligns with users’ practical needs, rather than develop LLM agents that directly automate tasks.
@@ 3.1 Task Creation
We rely on LLMs to generate the tasks. To fully exploit the functionalities, we adopted two strategies to guide LLMs to “search” for possible tasks, i.e., a top-down approach that starts from high-level functionalities of the software, followed by a bottom-up approach that considers low-level APIs supported by the software.
@@
A unique opportunity and challenge in the Software scripting interface setting is that we have a list of all APIs supported by the software, denoted as $\mathcal{A}=\{a_{1},a_{2},...,a_{n}\}$, and the scripts largely rely on the “collaboration” of multiple APIs. Therefore, we prompt LLMs with appropriate API combinations to spark LLMs’ knowledge about the software. Given the large number of APIs and the potential for multiple tasks to be accomplished using the same set of APIs, we introduce an API synergy graph, $\mathcal{G}$, to model the likelihood of APIs working together. Graph-based representations are widely employed to model relationships in related applications such as API recommendations Qi et al. (2022); Huang et al. (2022) LLM-based reasoning Anokhin et al. (2024); Liu et al. (2024a), and so on Wu et al. (2022); Xu et al. (2025).
@@ 3.2 Skill Generation with Trials
One advantage of generating skills offline is that we can generate the code with multiple trials offline without adding any burden to the users during runtime, providing a better user experience. For each generated $t_{i}$, we use a strategy similar to the ideas of Wang et al. (2024b, a), where the LLM learns from the execution feedback in the software to refine its own outputs. Additionally, we use another LLM or a Large Vision Language Model (LVLM) as a validator to judge the script by looking at the generated code, execution output, and (visual) outcome in the software. The validator comments on whether the skill accomplishes the task and provides feedback for improvement for the next trial. We show the system and user prompt for the validator in Table 8. A skill is added to the skillset $\mathcal{S}$ if it passes the validator. Specifically, when refining the scripts, the LLM receives a structured prompt containing the task description, code from the previous round, any execution errors generated by the software, and the validator’s assessment (including suggestions for improvement). This structured feedback enables the LLM to refine the script by addressing both code-level issues and misalignments with task intent. We allow up to three trials per task. We show the prompt for ExtendScript code generation in Table 7. Example skills with above-mentioned elements are shown in Appendix C.
@@ 4 Experiments
Adobe Illustrator is a leading vector graphics software used by professionals worldwide for tasks such as creating logos, illustrations, and complex design elements. It supports multiple categories of high-level functionalities, including drawing, arranging objects, and applying effects. A full list of the high-level functionalities used in the top-down search in Section 3.1 is provided in Appendix A, which can be obtained through Illustrator’s official scripting guide.  [^3] https://helpx.adobe.com/pdf/illustrator_reference.pdf Modern software platforms that support scripting interfaces offer tasks with varying levels of complexity. This study focuses on generating “atomic” skills: fundamental, modular tasks that involve minimal design choices and serve as building blocks for more complex operations. For instance, in the context of Illustrator, instead of creating an entire flower design, an atomic skill would arrange pre-designed petals into a circular pattern, emphasizing precision and modularity. Illustrator’s scripting interface supports $1818$ API endpoints, where $378$ are the methods and the remaining are attributes of the object, enabling extensive programmatic control over the software. We also explored using Excel as the testbed, but were unable to conduct a large-scale experiment due to limitations in Excel’s ability to permit programmatic control over Office Script from outside applications. See Appendix E for details.
@@
Adobe Illustrator, as a vector design tool, focuses on creating and manipulating graphical objects. In practical applications, a typical Illustrator project often contains numerous objects, and scripts are usually expected to operate on specific subsets of these objects. The common approach is to manually select the desired objects before running the script, leveraging the selection attribute available for each object. Therefore, for each skill, we prompt the LLM to generate initialization scripts that set up the document with necessary elements and adjust their selection attributes according to the task description. Following previous work with other virtual environments Wang et al. (2024b), we prompt the LLM to make the task code generic and reusable, as well as other Illustrator-specific instructions. The full prompt is shown in Appendix B. Table 1 presents an example skill, including the corresponding task description, code implementation, and the layouts after running the initialization script and the skill, respectively.
We now detail how we operationalize our framework in Adobe Illustrator. We first execute the top-down simulation using the functionality categories listed in Appendix A, where the LLM explores tasks in each subcategory over three rounds. The validated skills, combined with sample scripts provided by the software, are then used to construct the API synergy graph. To train the GCN model for link prediction, we randomly split the sampled edges (both positive and negative) into training, development, and test sets with a ratio of 0.85/0.05/0.1. The GCN model has two layers with an embedding size of $128$. We trained the model for $300$ epochs using the Adam optimizer with a learning rate of $0.01$. Next, for each of the approximately 500 method APIs, we retrieve its top-$k$ ($k=5$) synergistic APIs and prompt the LLM to generate related tasks for one round. [^4] We selected $k=5$ to balance between capturing potentially synergistic APIs and avoiding overwhelming the LLM with irrelevant API information. In each round of the task creation stage (both top-down and bottom-up approaches), the LLM generates ten tasks. For each task, the LLM is allowed up to three trials with previous feedback to generate the code of skills. We note that these hyperparameters may be subject to change based on cost considerations. We chose these values to ensure that the cost and size of the generated skillset remain manageable.
@@
To curate a test set that represents the needs of real users, we iteratively prompt the LLM to generate useful tasks in Illustrator. We manually verify the tasks and correct any issues in the corresponding initialization scripts, if any, to enable automatic evaluation, resulting in $94$ test tasks. We use the all-mpnet-base-v2 model from sentence BERT Reimers and Gurevych (2019) to encode the task descriptions for both test set and the skillset built during offline simulation. We use the test task description as the query and retrieve the relevant skills from the skillset with semantic matching through cosine similarity.
@@
The baseline method for this task is runtime generation where an LLM generates code to solve the query during runtime. We use gpt-4o 2024-08-06 as the LLM for baseline, as well as the ones in the proposed framework for fair comparisons. We note that advances in program synthesis methods are complementary to our approach; a stronger program synthesis model would benefit both the baseline and our framework. To demonstrate the generalization ability of our framework, we also evaluated lama-3.1-70B and deepseek-r1 as the code generation model on a random sample of $200$ tasks from each of the top-down and bottom-up strategies.
Proposed Methods
(1) Retrieval-Only (RO): Given a user query $q$, the system retrieves the most semantically relevant skill using cosine similarity over sentence embeddings. This approach ensures low-latency responses and minimal runtime cost, and avoids unsafe outputs, as all retrieved skills are pre-validated through offline trials. (2) Retrieval-Augmented Generation (RAG): To explore the use of the skillset for handling arbitrary user queries, we prompt an LLM with the top-$r$ ($r=3$) retrieved skills as in-context examples. This enables flexible adaptation and composition beyond atomic tasks.
@@ 5.1 Effectiveness in Real-World Scenarios
Our proposed method demonstrates a substantial improvement over the baseline in automating tasks within Adobe Illustrator’s scripting environment. As shown in Table 2, our approach (RO) achieves a success rate of $44.7\%$, outperforming the baseline’s success rate of $28.7\%$ on the held-out test set. This evaluation setup effectively measures the end-to-end performance of automation systems when deployed in real-world software contexts, with the natural language task descriptions in the test set simulating real user queries. In addition to RO, the RAG variant achieves a success rate of $42.6\%$, improving baseline performance by approximately $15\%$. This demonstrates that retrieved skills not only serve as executable scripts but also act as effective in-context examples for code generation.
A key advantage of RO lies in its markedly lower runtime costs, which is a critical consideration given the potential volume and repetitive nature of user queries. RO avoids runtime code generation entirely by retrieving pre-validated skills, resulting in an average response time of just $0.1$ seconds for retrieval and zero token cost at runtime, as offline simulation is a one-time expense. In contrast, both the baseline and RAG require significantly more resources at runtime, with response times of $4.0$ and $4.3$ seconds, and token costs of $666$ and $1219$.
@@
RAG 42.6% 4.3 s 1219
RO 44.7% 0.1 s 0*
Table 2: Evaluation of the baseline and the proposed approaches. * The token cost for offline simulation in our approach is one-off and not included in the runtime cost.
@@ 5.2 Effectiveness of Two Simulation Strategies
This outcome is expected, given the complexity and long-tail nature of software scripting API design. For example, a suite of APIs in Illustrator relates to physically printing, exporting to various file formats, and so on. These are “end-of-workflow” commands or meta-actions about the environment, rather than modular skills for manipulating document content. Verifying their success is difficult in a closed-loop simulation. Some APIs are tied to specific legacy software versions. Therefore, complete API coverage is an impractical goal. However, the long tail still represents numerous specialized APIs that are infrequently used and therefore less likely to be well-represented in the LLMs’ training data. This highlights room for future improvement, which we discuss in Section Limitations.
@@
Test Set Top-down Bottom-up
Avg. Score 2.48 2.28 1.75
Table 5: Average usefulness scores (1–3 scale) of sampled tasks, as rated by human evaluators.
Human-perceived usefulness
To assess the practical value of the discovered skills, we conducted a small-scale user study in which two Illustrator users with programming experience rated the usefulness of 150 sampled tasks on a 3-point scale: 1 (very rarely useful), 2 (occasionally helpful), and 3 (very useful). The sample included 50 tasks each from the test set, top-down skills, and bottom-up skills. Annotators judged each task based solely on its description. We discuss the annotation setup and annotator agreement in Appendix D. As shown in Table 5, skills produced by the top-down simulation strategy received a relatively high average usefulness score of $2.28$, closely matching the $2.48$ average for tasks in the test set. In contrast, bottom-up skills averaged $1.75$. Although this is relatively lower, it still leans toward the “occasionally helpful” category. This result is consistent with the bottom-up strategy’s goal of uncovering long-tail functionalities. We examine these differences further in the qualitative analysis below.
@@
The top-down simulation tends to generate tasks that reflect high-level design intentions familiar to users. For instance, it produces skills such as “arranging selected objects in a circle” (Table 1), along with other layout patterns like zigzag or starburst formations. These tasks are aligned with common design workflows and are typically easy to interpret and apply. The bottom-up simulation yields more technically specialized tasks that leverage less frequent APIs or scripting features. One example (Table LABEL:tab:full_example_2) involves alerting the user when selected objects have tags that match a given list. Such skill introduces conditional logic and metadata inspection, which is expected to be rarer in practical usage but provides greater flexibility for advanced users.
@@
During bottom-up simulation, we use a GNN-based link prediction model to model the synergy of APIs. We show that this model can find synergistic APIs more effectively, compared to semantically matching similar APIs. We evaluate the performance of the GCN model in retrieving synergistic APIs for task creation using the Hit@$k$ metric. In this context, Hit@$k$ evaluates, for each API, how often the correct items are within the top-$k$ predicted items where correct items refer to APIs in existing skills (not in the training set). We compare the link prediction model with a semantic matching baseline, which retrieves top-$k$ similar APIs using all-mpnet-base-v2 embeddings and cosine similarity Table 6 shows our model significantly outperforms semantic matching in identifying synergistic APIs, enhancing task creation in bottom-up simulation.
@@
One key advantage of our offline simulation is the ability to perform multiple trials during code generation without incurring runtime penalties. We analyzed the impact of allowing up to three trials for code generation on the overall success rate. Table 7 demonstrates that the success rate nearly doubles after three trials for both top-down and bottom-up strategies. Specifically, the success rate for both top-down and bottom-up strategies nearly doubles after three trials. We found similar improvement when using Llama-3.1-70B and deepseek-r1 as the code generation models shown in Table 11.
@@ 5.3 Reliability of LVLM’s judgment
We employed an LVLM to evaluate whether the execution outcomes satisfy the task descriptions. To estimate the reliability of the LVLM’s judgments, we compared its assessments with human evaluations on a sample of $122$ tasks. Table 8 shows that the LVLM has a precision of $90.9\%$ and a recall of $80.0\%$. Out of $50$ tasks deemed successful by humans, the LVLM correctly identified $40$ as successful and $10$ as failures. Meanwhile, the LVLM only incorrectly labeled $4$ tasks as successful when they were not (false positives). These results indicate that the LVLM is generally reliable, albeit slightly conservative in its judgments.